OpenStudy (anonymous):

A boat which moves at 5m/s in still water crosses a river without drift in 2min if the width of the river is 480m then the river flows with a a speed of?

OpenStudy (callisto):

I'm not sure if it is correct...  Let x be the speed of river flow. Resultant speed = 480/(2x60) = 480 / 120 = 4ms^-1 So, 5+x = 4 x = ...